OpenStudy (anonymous):

What is the exact area of a circle with a radius of 12 feet? A. 24π square feet B. 36π square feet C. 121π square feet D. 144π square feet

OpenStudy (anonymous):

Area = PI * R^2 R = 12, so R^2 = 144 A = 144 * PI.

OpenStudy (anonymous):

A=πr2. So, A=π×122. That's 144πfeet2.

OpenStudy (anonymous):

CAN YOU HELP WTH 4 MORE sorry caps

OpenStudy (anonymous):

ya sure i can try

OpenStudy (anonymous):

i got 75.12 but that isnt an answer choice so i cant really help you

OpenStudy (anonymous):

i did this in school an i passed

OpenStudy (anonymous):

*and

OpenStudy (anonymous):

What is the exact area of a circle with a diameter of m? A. 25m² B. 50m² C. 156.25m² D. 625m²

OpenStudy (anonymous):

of 25m

OpenStudy (anonymous):

D

OpenStudy (anonymous):

What is the approximate area of a circle with a radius of ft? A. 81.6 square feet B. 163.3square feet c. 530.7 square feet D. 2122.6 square feet

OpenStudy (anonymous):

of 26ft

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@adriwolf are u going to explain why it is that answer

OpenStudy (anonymous):

26 * 26 = 676 and 676 * 3.14 = 2122.6

OpenStudy (anonymous):

What is the approximate area of a circle with a diameter of 30 inches? Use 3.14 A. 94.2 square inches B. 188.4 square inches C. 706.5 square inches D. 2826 square inches

OpenStudy (anonymous):

30/2 = 15 and 15 * 15 = 225 * 3.13 = 706.5

OpenStudy (anonymous):

its not d.?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

What is the approximate area of the circle? Use 3.14 A. 17.6 square inches B. 24.6 square inches C. 35.2 square inches D. 98.5 square inches

OpenStudy (anonymous):

you need to multiply the radius not the diameter

OpenStudy (anonymous):

5.6 * 5.6 = 31.36 and 31.63 * 3.14 = 98.5

OpenStudy (anonymous):

80% THX